• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name cisbani

    Log entry time 00:04:08 on November13,2008

    Entry number 248474

    keyword=Swing shift summary

    Swing Shift Summary (Nov/12/2008)

    16:00 Ask MCC to move BPMB x at -0.5 and raster size 2x2 MCC (2 uA current, on 3He)

    16:30 Target Calibration (EPR measurement)

    16:45-20:45 Moller measurement (see entry 248427)

    20:45 Do a raster scan to check that the beam is centered
    Raster T1 T3 T5 Spot
    3x3MCC 17.9k 253 1.50 2x1.5 mm
    4x4MCC 18.5k 273 1.47 3x2.5 mm
    5x5MCC 17.6k 273 1.47 3.5x3 mm
    6x6MCC 17.6k 256 1.65 4x3 mm
    6.58x6.58MCC 16.4k 245 1.42 5.5x4.5 mm

    21:30 Two RICH pedestal runs 3302 and 3303

    21:45 Production run 3304 (no Happex yet), (T1 and T3 at 300 Hz for calibration), beam current at 8 uA, Target 3He transverse pol. 62 %

    22:15 Happex Run started 30403, production run 3305 (T3 prescaled to get 50 Hz)
    Run stopped to do HV adjustment

    22:22 After the BB chamber tuning, finally started the production run 3306

    23:05 Lost beam due to RF failure; 15-20 minutes to get it back; run 3307 stopped

    23:20 Beam back, started new production run 3308